# Enhancing Wide-Angle Image Using Narrow-Angle View of the Same Scene

### v2

@@ 4.5 Enhanced image reconstruction pipeline
4.6 Training Details
Key hyperparameters for training and evaluation included a patch size of 128 × 128, Adam optimizer with a learning rate of 0.0001 with exponential decay every 10 epochs, a batch size of 16, perceptual loss layers (Conv2_2, Conv3_2, Conv4_2 of VGG19), and a threshold for patch similarity empirically set to 0.7.
The generator network was pre-trained for 50 epochs using content and visual losses before incorporating adversarial loss, each epoch carrying 10 sample images. The full generator-discriminator framework was then trained for an additional 2000 epochs, totaling 20.5K iterations.
The convergence behavior of the generator loss has been illustrated in Figure 4. These curves represent the first 15 epochs where the generator has been trained solo. Upper row features the raw loss curves as generated by the GAN architecture, while the lower row has been smoothened using the numpy convolve [33] method.
Narrow FoV Input wide FoV Our result
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 2: Experimental results showing the output generated by our model when processing narrow and wide input images.
[H] Dataset bicubic Multiscale GAN Ours Set5 0.5657 0.2823 0.0795 Set14 0.6114 0.3717 0.2032 BSD100 0.5726 0.4112 0.3361
@@ 4.6 Training Details
Figure 6: Visual comparison of our results against the mentioned prior arts
@@ 4.7 Experiment Results
Table 2 presents the results of our experiment, where a narrow and a low-resolution wide image are used as input. These images are processed through our model to generate the final output.
Table 1 illustrates the cascading effect of the outputs derived from images captured at different zoom levels. In the first row, the input images consist of a 2x (telephoto) and a 1x (standard) zoom, which collectively produce a high-resolution 1x (standard) output. The second row takes the output from the previous row as one input and combines it with a low-resolution 0.5x (wide) image, ultimately generating a high-resolution 0.5x (wide) image. This sequence effectively simulates how our method can be applied in a cascading manner using all available zoom lenses.
@@ 5 Comparison with Prior Arts
We have evaluated our proposed methods on PSNR, SSIM and LPIPS [16] metrics over Set5 [34], Set14 [35] and BSD100 [36] datasets. For PSNR and SSIM, we have compared our results with existing technologies such as ESRGAN [7], LatticeNet [12], ELAN [13], DLGSANet [14] and DMFFN [15] in Table 4 and Table 5. We have presented our comparison with bicubic interpolation and Multiscale GAN [37] on the LPIPS metric in Table 3. Visual comparison of our method against prior super-resolution models has been presented in Figure 6.
@@ 6 Conclusion
In this work, we presented a novel approach for enhancing ultra-wide field-of-view (FoV) images by leveraging paired narrow-FoV views of the same scene. The core of our method is a patch-based cross-view attention mechanism that selectively transfers fine-grained details from narrow-FoV patches to their wide-FoV counterparts, guided by Gram matrix-based visual encoding that captures texture statistics for accurate alignment. To further refine the reconstruction, we introduce a cascading lens stack mechanism that progressively transfers details across adjacent lens images, facilitating spatial consistency and coherent integration across the full image.
We trained our model on the Landscape HQ dataset to ensure coverage of both structural diversity and fine detail. For evaluation, we used standard benchmark datasets including BSD100, Set5, and Set14. Experimental results demonstrate that our method consistently outperforms strong baselines such as ESRGAN, ELAN, DLGSANet, DMFFN, and LatticeNet in both PSNR and SSIM metrics. Qualitative comparisons further confirm the perceptual superiority of our results.
Overall, our framework successfully addresses the challenge of enhancing ultra-wide FoV images by combining cross-view attention, texture-aware encoding, and a cascading lens-based reconstruction strategy. Future extensions could explore its application to video enhancement and real-time inference scenarios.

### v3

@@ 4.3 Image Reconstruction Pipeline
Figure 6: Visual comparison of our outcomes with existing methods.
Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
RealSRv3 0.2940 0.2907 0.2582 0.3409 0.3054 0.1740
ImagePairs 0.2979 0.2775 0.2754 0.2866 0.2872 0.2435
@@
4.4 Training Details
Key hyperparameters for training and evaluation included an Adam optimizer with a learning rate of 0.0001 with exponential decay every 10 epochs, a batch size of 16, perceptual loss layers (Conv2_2, Conv3_2, Conv4_2 of VGG19), and a threshold for patch similarity empirically set to 0.7. During training, an image is partitioned into 64 patches, and the patch dimensions are determined likewise, for example $128\times 128$ for Landscape dataset images. In case of inference, image dimensions are resized to the next multiple of 64, if needed, to facilitate the partitioning strategy.
The generator network was pre-trained for 50 epochs using content and visual losses before incorporating adversarial loss, each epoch carrying 10 sample images. The full generator-discriminator framework was then trained for an additional 2000 epochs, totaling 20.5K iterations. The convergence behavior of the generator loss has been illustrated in Fig 4. These curves represent the first 15 epochs where the generator has been trained solo. Upper row features the raw loss curves as generated by the GAN architecture, while the lower row has been smoothened using the numpy convolve [25] method.
4.5 Experimental Setting
For evaluation against existing works in the field, we have used 2 benchmark datasets which are RealSRv3 [26] and ImagePairs [27]. These evaluation datasets differ from classical SR datasets in that the narrow and wide-FoV pairs are captured using actual focal length manipulation, as opposed to mathematical simulation, and hence are better suited to our particular case. We have, however, performed a parallel validation on traditional datasets Set5 [28], Set14 [29], and BSD100 [30], results of which are available in the supplementary document (Section 4). All reported metrics have been calculated based on the complete output image, as opposed to patch-wise measures. We have used the LR image provided by the dataset as the input, and the given HR as the ground truth.
@@ 4.6 Experiment Results
Table 1 illustrates the cascading effect of the outputs derived from images captured at different zoom levels. In the first row, the input images consist of a 2x (telephoto) and a 1x (standard) zoom, which collectively produce a high-resolution 1x (standard) output. The second row takes the output from the previous row as one input and combines it with a low-resolution 0.5x (wide) image, ultimately generating a high-resolution 0.5x (wide) image. This sequence effectively simulates how our method can be applied in a cascading manner using all available zoom lenses. Visual outcomes on the benchmark datasets have been elaborated in the supplementary document (Section 3B).
Loss convergence in the epoch interval of 5 to 10 has been shown in Fig 4. The lower curve has been smoothed to denote the overall downward trend of the graph. Details on loss minimization has been included in the supplementary document (Section 3A).
@@ 5 Comparison with Prior Arts
We have evaluated our proposed methods on LPIPS [31], PSNR, and SSIM metrics over datasets mentioned earlier. For all the metrics, we have compared our results with existing technologies such as ESRGAN [7], LatticeNet [12], ELAN [13], DLGSANet [14] and DMFFN [15] and presented the results in Tables 4, 4, and 4 respectively. We have also included a visual comparison against the prior methods in the Fig 6.
5.1 Quantitative Analysis
In terms of perceptual quality, as assessed using the LPIPS metric, our approach shows a clear improvement over earlier methods. This is largely due to the integration of visually similar patches, the selection of representative perceptual features, and the inclusion of perceptual loss within the discriminator, all of which contribute to more visually realistic outputs.
Similarly, our model achieves a noticeably higher PSNR score compared to other techniques. Notably, on the ImagePairs dataset, it is the only method to surpass a PSNR value of 30.
In the case of SSIM, our method also achieves a marked advantage on the RealSRv3 dataset, indicating superior preservation of structural details. However, for the ImagePairs dataset, the DMFFN model slightly outperforms ours. This can likely be attributed to its Multi-stage Feature Supplementation, which enhances geometrical consistency by combining shallow and deep features—albeit with increased computational demands.
5.2 Qualitative Analysis
We evaluated 2× super-resolved outputs from the referenced datasets against those produced by existing models. First, many of these models lack broader content-aware guidance, making it difficult for them to accurately reconstruct structural details—this is especially evident in elements like building windows (first image). Second, earlier methods tend to produce less distinct edges, often resulting in soft or blurred boundaries, whereas our approach delivers noticeably sharper edges, as illustrated by the leaf contours in the second image. Lastly, in visually rich scenes with high color saturation and contrast—such as green grass (third image)—previous models often fail to preserve vibrancy and fine texture. In contrast, our method successfully maintains these visual attributes by utilizing cues from the narrow field-of-view input, which is rich in detail.
@@ 6 Conclusion
This paper introduces an approach for enhancing ultra-wide field-of-view (FoV) images by leveraging paired narrow-FoV views of the same scene. The core of our method is a patch-based cross-view attention mechanism that selectively synthesizes fine-grained details from narrow-FoV patches into their wide-FoV counterparts, guided by a Gram matrix-based visual encoding to capture the texture statistics for accurate alignment. To further refine the reconstruction, we introduce a cascading lens stack mechanism that progressively transfers details across adjacent lens images, facilitating spatial consistency and coherent integration across the full image.
The model trained is on the Landscape HQ dataset to cover both structural diversity and finer details, and fine-tuned on DIV2K for handling rectangular images. For evaluation, we have used standard benchmark datasets including RealSRv3 and ImagePairs. Experimental results demonstrate a promising improvement over strong baselines such as ELAN, DLGSANet, DMFFN, and LatticeNet, especially in LPIPS, along with both PSNR and SSIM metrics. Qualitative comparisons further add to the perceptual superiority of our results.
Limitations to our current work include the necessity of further fine-tuning our model with real-life ultra-wide and narrow FoV pairs, since it is trained on synthetic data. Also, additional optimizations are required to deploy the model on-device to serve real-time purposes and video super-resolution tasks.
Declarations
  • Funding: This research did not receive external funding.
  • Competing Interests: The authors do not have any competing interests as defined by Springer, or other interests that might be perceived to influence the results and/or discussion reported in this paper.